# Causal Inference for a Hidden Treatment

### v2

@@ 3 Identification of ATE
$\begin{split}\delta(X,a^{*})=&E(Y\mid A=a^{*},X)-E(Y\mid A^{*}=a^{*},X)\\ =&\{E(Y\mid A^{*}=1-a^{*},X)-E(Y\mid A^{*}=a^{*},X)\}\textrm{pr}(A^{*}=1-a^{*}\mid A=a^{*},X),a^{*}=0,1.\end{split}$ (8)
This bias will generally be non-null, unless the ATE of $A^{*}$ on $Y$ is actually null, or if misclassification error is absent. Although unlikely, it may also happen that under certain exceptional laws, the conditional bias $\delta(X)$ given $X$ itself is not null, with opposite signs over different parts of the support of $X$, such that however it averages over $X$ to equal zero.
In order to identify $\psi_{a^{*}}$, we carefully leverage the information about $A^{*}$ contained in $Y$, the proxy $A$, and the surrogate $Z$. Formally, our hidden treatment model imposes Assumptions 2 and 3 below.
[…]
Figure 1: $A^{*}$ is the hidden binary exposure, $A$ is a binary surrogate, $Z$ is a binary proxy, $Y$ is the outcome, and $X$ is the confounder.
@@
We implicitly assume that $A\neq A^{*}$ and $Z\neq A^{*}$, a necessary implication of misclassification error. If $A$ equals $A^{*}$, the matrix $P(Y,Z,A=1)$ which will be introduced shortly, would not be of full rank. Therefore, a non-zero determinant of $P(Y,Z,A=1)$ indicates that $A\neq A^{*}$. A similar argument applies to $Z$. This re-iterates the fact that we are primarily concerned with the most challenging settings where misclassification error cannot be ruled out empirically.
Under Assumptions 2 and 3, the joint distribution of $(A^{*},Y,A,Z,X)$ is determined up to labeling. To illustrate the idea, consider a simplified scenario where $Y$ is binary and $X=\emptyset$. In this situation, the joint distribution of the observed data $(Y,A,Z)$ has $72^{3}-1=$ 7 degrees of freedom, and we have 7 unknown parameters: $E(A^{*}),E(Y\mid A^{*}=a^{*}),E(A\mid A^{*}=a^{*}),E(Z\mid A^{*}=a^{*}),a^{*}=0,1$. Let $P(Y,Z)=(\textrm{pr}(Y=2-i,Z=2-j))_{i,j}$, $P(Y,Z,A=1)=(\textrm{pr}(Y=2-i,Z=2-j,A=1))_{i,j}$, $P(Y\mid A^{*})=(\textrm{pr}(Y=2-i\mid A^{*}=2-j))_{i,j}$, $P(Z\mid A^{*})=(\textrm{pr}(Z=2-i\mid A^{*}=2-j))_{i,j}$, $P_{D}(A^{*})=\text{diag}(\textrm{pr}(A^{*}=1),\textrm{pr}(A^{*}=0))$, $P_{D}(A=1\mid A^{*})=\text{diag}(\textrm{pr}(A=1\mid A^{*}=1),\textrm{pr}(A=1\mid A^{*}=0))$. By Assumption 3, we have
$P(Y,Z)$ $=P(Y\mid A^{*})P_{D}(A^{*})P(Z\mid A^{*})^{\mathrm{\scriptscriptstyle T}},$
$P(Y,Z,A=1)$ $=P(Y\mid A^{*})P_{D}(A^{*})P_{D}(A=1\mid A^{*})P(Z\mid A^{*})^{\mathrm{\scriptscriptstyle T}}.$
By Assumption 2, $P(A\mid Z)$ is invertible. Multiplying $P(Y,Z,A=1)$ by $P(A\mid Z)^{-1}$ gives
$P(Y,Z,A=1)P(Y,Z)^{-1}=P(Y\mid A^{*})P_{D}(A=1\mid A^{*})P(Y\mid A^{*})^{-1},$
which as noted in Hu (2008) is the canonical form of the eigendecomposition of the matrix on the left-hand side. Since the right-hand side is observed, we can solve for the eigenvalues and eigenvectors, which correspond to the diagonal elements of $P_{D}(A=1\mid A^{*})$ and the columns of $P(Y\mid A^{*})$, respectively. While eigenvectors are identified up to a scale, the requirement that columns of $P(Y\mid A^{*})$ sum to 1 fixes it. The remaining challenge lies in correctly matching each eigenvalue with its respective diagonal element in $P_{D}(A=1\mid A^{*})$. To resolve this, we propose the following assumption which encodes four alternative ways in which the surrogate $A$ carries sufficient information about $A^{*}$ to resolve the remaining identification gap.
@@
At least one of the following conditions holds: for any value $x$ of $X$,
@@ 4. Condition 4
Assumption 4 clarifies the different roles of $A$ and $Z$, although they appear symmetric in Figure 1. While both the proxy $Z$ and the surrogate $A$ need to satisfy Assumptions 2 and 3, an extra Assumption 4 is imposed on the surrogate. This assumption essentially suggests that $A$, as a surrogate, should provide more information about $A^{*}$ than a random guess. The condition can reasonably be expected to hold if $A$ is a useful measurement of $A^{*}$ in the sense that it reflects a genuine albeit imperfect effort to measure the latter. Condition 1 is equivalent to $\textrm{pr}(A=1\mid A^{*}=0,X=x)<\textrm{pr}(A=1\mid A^{*}=1,X=x)$. Similarly, Condition 2 can be written as $\textrm{pr}(A^{*}=1\mid A=0,X=x)<\textrm{pr}(A^{*}=1\mid A=1,X=x)$. It is noteworthy that Condition 1 and Condition 2 are effectively of similar but symmetric nature. Condition 3 indicates the precision $\textrm{pr}(A^{*}=1\mid A=1,X=x)>0.5$, while Condition 4 states that the false omission rate $\textrm{pr}(A^{*}=1\mid A=0,X=x)<0.5$. These conditions are intuitively reasonable in many scenarios. For example, consider $A^{*}$ as the actual COVID-19 vaccination status of a patient, and $A$ as whether the patient was vaccinated at a specific hospital. If hospital records show a positive vaccination status, it must be that the patient was truly vaccinated, making $\textrm{pr}(A^{*}=1\mid A=1,X=x)=1$, thereby fulfilling Condition 2 (and Condition 1) of Assumption 4. If one can reasonably assume that the misclassification error is balanced across covariates, such that $X\!\perp\!\!\!\perp A\mid A^{*}$, i.e., the surrogate $A$ only depends on $A^{*}$, then one can safely omit $X$ from Assumption 4. We note that Hu (2008) considers alternatives to Conditions 1-4.
The informal proof given above is formalized in Theorem 1 below in a more general setting where $Y$ can be either binary, a count, or a continuous outcome, and $X$ is possibly vector-valued. For the proof of this theorem, please refer to §LABEL:supp-sec:thm_identification in the Supplementary Materials.
@@
It is important to note that a key step in the proof of the theorem entails establishing point identification of the joint conditional oracle law $\textrm{pr}(A,Z,Y,A^{*}\mid X=x)=\textrm{pr}(A\mid A^{*},X=x)\textrm{pr}(Z\mid A^{*},X=x)\textrm{pr}(Y\mid A^{*},X=x)\textrm{pr}(A^{*}\mid X=x)$, which, in principle implies point identification of any smooth functional of the oracle law. This observation is crucial to establishing identification for more general causal functionals beyond the ATE. The next section builds on this key identification result to construct highly robust and efficient semiparametric estimators of $\psi_{a^{*}}$ under the semiparametric model $\mathcal{M}$ for the observed data, induced by the oracle semiparametric model $\mathcal{M^{*}}$ defined by the conjunction of Assumptions 1, 2, 3, and 4.
@@ 4 Semiparametric theory
Having established identification for $\psi_{a^{*}}$, our next objective is to characterize the tangent space of the semiparametric model $\mathcal{M}$ and to characterize its ortho-complement. By using some key results from modern semiparametric theory, we obtain a set of influence functions for $\psi_{a^{*}}$, which in turn motivates a feasible set of estimating equations. This step is crucial for developing efficient semiparametric estimators for $\psi_{a^{*}}$ and for conducting principled inferences under $\mathcal{M}$.
@@
$\Lambda^{F,\bot}=\left\{K=k\left(Y,A,A^{*},Z,X\right):E\left(K\mid Y,A^{*},X\right)=E\left(K\mid A,A^{*},X\right)=E\left(K\mid Z,A^{\ast},X\right)=0\right\}\cap L_{0}^{2,F},$ (10)
@@
$\begin{split}\Lambda^{F,\bot}=\Bigl{\{}W=M-E\left(M\mid Y,A^{*},X\right)-E\left(M\mid A,A^{*},X\right)-E\left(M\mid Z,A^{*},X\right)+2E\left(M\mid A^{\ast},X\right):\\ \qquad\qquad M=m\left(Y,A,A^{*},Z,X\right)\in L_{0}^{2,F}\Bigr{\}}.\end{split}$ (11)
The derivation of (10) and (11) is deferred to §LABEL:supp-sec:proof_tangentspace_full in the Supplementary Materials. In the following, we will make reference to the oracle nonparametric model $\mathcal{M}_{np}^{*}$ as the collection of oracle laws compatible with the DAG given in Figure 2 which places no restriction on the joint distribution of $(A^{*},Y,Z,A,X)$. We then have the following result:
@@
The efficient influence function of $\psi_{a^{*}}$ in the oracle hidden treatment model $\mathcal{M^{*}}$ is equal to the efficient influence function in the oracle nonparametric model $\mathcal{M}_{np}^{*}$.
@@
[…]
Figure 2: A DAG for the oracle nonparametric model $\mathcal{M^{*}}_{np}$.
The efficient influence function for $\psi_{a^{*}}$ in $\mathcal{M^{*}}_{np}$, denoted by $\phi_{a^{*}}^{F,eff}$ is well-established, and is equal to (2). Therefore, the set of influence functions for $\psi_{a^{*}}$ in the full data model $\mathcal{M^{*}}$ is the linear variety $\phi_{a^{*}}^{F,eff}\oplus\Lambda^{F,\bot}$ (Tsiatis, 2007).
So far, we have characterized the tangent space, the set of influence functions and the efficient influence function for $\psi_{a^{*}}$ in $\mathcal{M^{*}}$. We now turn our attention to the observed data model $\mathcal{M}$. The following result summarizes our characterization of the tangent space of the observed data hidden treatment model $\Lambda$, and its ortho-complement, obtained by carefully applying a certain coarsening operator which maps the tangent space for the full data model, $\Lambda^{F}$, onto the $L^{2}_{0}(Y,A,Z,X)$ space.
@@
$\begin{split}&\Lambda^{\bot}=\bigl{\{}\{A-E(A\mid A^{*}=0,X)\}\{Z-E(Z\mid A^{*}=1,X)\}h_{1}(Y,X)+\{A-E(A\mid A^{*}=1,X)\}\\ &\qquad\quad\,\{Z-E(Z\mid A^{*}=0,X)\}h_{2}(Y,X):E\{h_{1}(Y,X)\mid A^{*},X\}=E\{h_{2}(Y,X)\mid A^{*},X\}=0\bigr{\}}\cap L^{2}_{0}.\end{split}$ (12)
@@
where $M_{Y\mid A^{*},X}$ is defined as
$\begin{bmatrix}f(Y=1\mid A^{*}=0,X)&f(Y=1\mid A^{*}=1,X)\\ \vdots&\vdots\\ f(Y=k\mid A^{*}=0,X)&f(Y=k\mid A^{*}=1,X)\end{bmatrix}.$
This implies that $h_{1}$ and $h_{2}$ belong to the null space of the $k\times 2$ matrix $M_{Y\mid A^{*},X}$. Under Assumption 2, the null space has a dimension of $k-2$ when $k\geq 2$. When $Y$ is binary, i.e., $k=2$, $M_{Y\mid A^{*},X}$ is full rank, leading to the unique solution of (13): $h_{j}(Y,X)=0,j=1,2$. This is further detailed in Corollary 1.
Corollay 1.
@@
$\begin{split}&\Lambda^{\bot}=\bigl{\{}\{A-E(A\mid A^{*}=0,X)\}\{Z-E(Z\mid A^{*}=1,X)\}w_{1}(Y,X)+\{A-E(A\mid A^{*}=1,X)\}\\ &\qquad\quad\,\{Z-E(Z\mid A^{*}=0,X)\}w_{2}(Y,X):w_{1},w_{2}\in P_{2}\bigr{\}}.\end{split}$ (14)
Such equivalence provides a closed-form characterization of $P_{1}$. Based on this, we obtain a closed-form characterization of $\Lambda^{\bot}$. See §LABEL:supp-sec:orthocomplement in the Supplementary Materials for a proof of the result.
@@
The following equation provides a closed form expression of an influence function for $\psi_{a^{*}}$ in the hidden treatment model $\mathcal{M}$ which applies whether $Y$ is binary, categorical, a count or a continuous variable:
@@
$\{Y-E(Y\mid a^{*},X)\}+E(Y\mid a^{*},X)-\psi_{a^{*}}.$ (16)
Equation (4) admits the triple robustness property: $E(\phi_{a^{*}})=0$ if one of the following holds.
  1. 1.
    $f(A\mid A^{*},X)$ and $f(Y\mid A^{*},X)$ are correctly specified.
  2. 2.
    $f(Z\mid A^{*},X)$ and $f(Y\mid A^{*},X)$ are correctly specified.
  3. 3.
    $f(A\mid A^{*},X)$, $f(Z\mid A^{*},X)$ and $f(A^{*}\mid X)$ are correctly specified.
The efficient influence function $\phi_{a^{*}}^{eff}$ is given by the orthogonal projection of $\phi_{a^{*}}$ onto the observed data tangent space of the hidden treatment model, i.e., $\phi_{a^{*}}^{eff}=\Pi(\phi_{a^{*}}\mid\Lambda)$ where $\Pi$ is the orthogonal projection operator. Therefore,
@@
In the case of binary $Y$, $w^{opt}_{1}=w^{opt}_{2}=0$, and thus (4) is the efficient influence function for $\psi_{a^{*}}$ in the hidden treatment model $\mathcal{M}$. A closed-form expression for $\{w^{opt}_{j}:j=1,2\}$ in the more general categorical outcome setting is given in the Supplemental Materials §LABEL:supp-sec:closed_form_w. As a closed-form expression for $\{w^{opt}_{j}:j=1,2\}$ and thus for the efficient influence function is not generally available for continuous $Y$, we give in the same Supplemental Materials Section a simple approach for obtaining an approximate efficient influence function based on a basis function approximation of $\{w^{opt}_{j}:j=1,2\}$.
The theorem provides an interesting insight about the relationship between an oracle influence function and the corresponding observed data influence function. Specifically, comparing (4) with the full data influence function (2), one may note that the former is obtained by substituting the indicator function $\mathbb{I}_{a^{*}}(A^{*})$ in the latter with the following expression:
$\frac{A-E(A\mid 1-a^{*},X)}{E(A\mid a^{*},X)-E(A\mid 1-a^{*},X)}\frac{Z-E(Z\mid 1-a^{*},X)}{E(Z\mid a^{*},X)-E(Z\mid 1-a^{*},X)}.$ (17)
Interestingly, (17) is an conditionally unbiased estimator of $\mathbb{I}_{a^{*}}(A^{*})$, in the sense that its conditional mean given $A^{*}$ and $X$ matches the indicator function $\mathbb{I}_{a^{*}}(A^{*})$:
$E\left\{\frac{A-E(A\mid A^{*}=1-a^{*},X)}{E(A\mid A^{*}=a^{*},X)-E(A\mid A^{*}=1-a^{*},X)}\frac{Z-E(Z\mid A^{*}=1-a^{*},X)}{E(Z\mid A^{*}=a^{*},X)-E(Z\mid A^{*}=1-a^{*},X)}\Bigm{|}A^{*},X\right\}$
$=$ $E\left\{\frac{A-E(A\mid 1-a^{*},X)}{E(A\mid a^{*},X)-E(A\mid 1-a^{*},X)}\Bigm{|}A^{*},X\right\}E\left\{\frac{Z-E(Z\mid 1-a^{*},X)}{E(Z\mid a^{*},X)-E(Z\mid 1-a^{*},X)}\Bigm{|}A^{*},X\right\}$
$=$ $\frac{E(A\mid A^{*},X)-E(A\mid 1-a^{*},X)}{E(A\mid a^{*},X)-E(A\mid 1-a^{*},X)}\frac{E(Z\mid A^{*},X)-E(Z\mid 1-a^{*},X)}{E(Z\mid a^{*},X)-E(Z\mid 1-a^{*},X)}$
$=$ $\mathbb{I}_{a^{*}}(A^{*}),$
where we have used the fact that $(A,Y,Z)$ are mutually independent given $A^{*},X$ as well as the fact that
$\frac{A-E(A\mid A^{*}=1-a^{*},X)}{E(A\mid A^{*}=a^{*},X)-E(A\mid A^{*}=1-a^{*},X)}$
and
$\frac{Z-E(Z\mid A^{*}=1-a^{*},X)}{E(Z\mid A^{*}=a^{*},X)-E(Z\mid A^{*}=1-a^{*},X)}$
are conditionally unbiased for $\mathbb{I}_{a^{*}}(A^{*})$. This provides an elegant link between (2) and (4). In the full data influence function (2), the unobserved variable $A^{*}$ only appears within the indicator function $\mathbb{I}_{a^{*}}(A^{*})$. In the observed data influence function (4), $\mathbb{I}_{a^{*}}(A^{*})$ is replaced by an expression of observed variables (17), whose expectaction conditional on $(A^{*},X)$, is equal to $\mathbb{I}_{a^{*}}(A^{*})$.
This observation suggests a general approach for deriving influence functions which is easily extended to other causal functionals of $(A^{*},X,Y)$ discussed in §2 under the hidden treatment model. We refer the reader to §6 for a detailed exposition.

### v3

@@ 3 Identification of ATE
In order to identify $\psi_{a^{*}}$, our hidden treatment model imposes Assumptions 2 and 3 below.
@@
[…]
Figure 1: (a) $A^{*}$ is the hidden binary exposure, $A$ is a binary surrogate, $Z$ is a binary proxy, $Y$ is the outcome, and $X$ is the confounder. (b) A DAG for the oracle nonparametric model $\mathcal{M^{*}}_{np}$.
[…]
@@
We implicitly assume that $A\neq A^{*}$ and $Z\neq A^{*}$, a necessary implication of misclassification error. If $A$ equals $A^{*}$, the matrix $P(Y,Z,A=1)$ which will be introduced shortly, would not be of full rank. Therefore, a non-zero determinant of $P(Y,Z,A=1)$ indicates that $A\neq A^{*}$. A similar argument applies to $Z$.
Under Assumptions 2 and 3, the joint distribution of $(A^{*},Y,A,Z,X)$ is determined up to labeling. To illustrate the idea, consider a simplified scenario where $Y$ is binary and $X=\emptyset$. In this situation, the joint distribution of the observed data $(Y,A,Z)$ has $72^{3}-1=$ 7 degrees of freedom, and we have 7 unknown parameters: $E(A^{*}),E(Y\mid A^{*}=a^{*}),E(A\mid A^{*}=a^{*}),E(Z\mid A^{*}=a^{*}),a^{*}=0,1$. Let $P(Y,Z)=(\textrm{pr}(Y=2-i,Z=2-j))_{i,j}$, $P(Y,Z,A=1)=(\textrm{pr}(Y=2-i,Z=2-j,A=1))_{i,j}$, $P(Y\mid A^{*})=(\textrm{pr}(Y=2-i\mid A^{*}=2-j))_{i,j}$, $P(Z\mid A^{*})=(\textrm{pr}(Z=2-i\mid A^{*}=2-j))_{i,j}$, $P_{D}(A^{*})=\text{diag}(\textrm{pr}(A^{*}=1),\textrm{pr}(A^{*}=0))$, $P_{D}(A=1\mid A^{*})=\text{diag}(\textrm{pr}(A=1\mid A^{*}=1),\textrm{pr}(A=1\mid A^{*}=0))$. By Assumption 3, we have $P(Y,Z)=P(Y\mid A^{*})P_{D}(A^{*})P(Z\mid A^{*})^{\mathrm{\scriptscriptstyle T}},P(Y,Z,A=1)=P(Y\mid A^{*})P_{D}(A^{*})P_{D}(A=1\mid A^{*})P(Z\mid A^{*})^{\mathrm{\scriptscriptstyle T}}.$ By Assumption 2, $P(A\mid Z)$ is invertible. Multiplying $P(Y,Z,A=1)$ by $P(A\mid Z)^{-1}$ gives $P(Y,Z,A=1)P(Y,Z)^{-1}=P(Y\mid A^{*})P_{D}(A=1\mid A^{*})P(Y\mid A^{*})^{-1},$ which as noted in Hu, (2008) is the canonical form of the eigendecomposition of the matrix on the left-hand side. Since the right-hand side is observed, we can solve for the eigenvalues and eigenvectors, which correspond to the diagonal elements of $P_{D}(A=1\mid A^{*})$ and the columns of $P(Y\mid A^{*})$, respectively. While eigenvectors are identified up to scale, the requirement that columns of $P(Y\mid A^{*})$ sum to 1 fixes it. The remaining challenge lies in correctly matching each eigenvalue with its respective diagonal element in $P_{D}(A=1\mid A^{*})$. To resolve this, we propose the following assumption encoding four alternative ways in which the surrogate $A$ carries sufficient information about $A^{*}$ to resolve the remaining identification gap.
@@
At least one of the following holds: for any value $x$ of $X$,
@@ 4. Condition 4
Assumption 4 clarifies the different roles of $A$ and $Z$, although they appear symmetric in Figure 1(a). While both the proxy $Z$ and the surrogate $A$ need to satisfy Assumptions 2 and 3, an extra Assumption 4 is imposed on the surrogate. This assumption essentially suggests that $A$, as a surrogate, should provide more information about $A^{*}$ than a random guess. The condition is expected to hold if $A$ is a meaningful but imperfect measurement of $A^{*}$. Condition 1 is equivalent to $\textrm{pr}(A=1\mid A^{*}=0,X=x)<\textrm{pr}(A=1\mid A^{*}=1,X=x)$. Similarly, Condition 2 can be written as $\textrm{pr}(A^{*}=1\mid A=0,X=x)<\textrm{pr}(A^{*}=1\mid A=1,X=x)$. Condition 1 and Condition 2 are effectively of similar but symmetric nature. Condition 3 indicates the precision $\textrm{pr}(A^{*}=1\mid A=1,X=x)>0.5$, while Condition 4 states that the false omission rate $\textrm{pr}(A^{*}=1\mid A=0,X=x)<0.5$. These conditions are reasonable in many scenarios. For example, consider $A^{*}$ as the actual COVID-19 vaccination status of a patient, and $A$ as whether the patient was vaccinated at a specific hospital. If hospital records show a positive vaccination status, it must be that the patient was truly vaccinated, making $\textrm{pr}(A^{*}=1\mid A=1,X=x)=1$, thereby fulfilling Condition 2 (and Condition 1). If one can assume that the misclassification error is balanced across covariates, such that $X\!\perp\!\!\!\perp A\mid A^{*}$, i.e., the surrogate $A$ only depends on $A^{*}$, then one can safely omit $X$ from Assumption 4. We note that Hu, (2008) considers alternatives to Conditions 1-4.
The informal proof given above is formalized in Theorem 1 below in a more general setting where $Y$ can be either binary, a count, or a continuous outcome, and $X$ is possibly vector-valued. For the proof of this theorem, please refer to §LABEL:supp-sec:thm_identification in the supplement.
@@
A key step in the proof entails establishing point identification of the joint conditional oracle law $\textrm{pr}(A,Z,Y,A^{*}\mid X=x)=\textrm{pr}(A\mid A^{*},X=x)\textrm{pr}(Z\mid A^{*},X=x)\textrm{pr}(Y\mid A^{*},X=x)\textrm{pr}(A^{*}\mid X=x)$, which, in principle implies point identification of any smooth functional of the oracle law. This observation is crucial to establishing identification for more general causal functionals beyond the ATE. The next section builds on this key identification result to construct highly robust and efficient semiparametric estimators of $\psi_{a^{*}}$ under the semiparametric model $\mathcal{M}$ for the observed data, induced by the oracle semiparametric model $\mathcal{M^{*}}$ defined by the conjunction of Assumptions 14.
@@ 4 Semiparametric theory
Having established identification for $\psi_{a^{*}}$, our next objective is to characterize the tangent space of the semiparametric model $\mathcal{M}$ and its ortho-complement. By using some key results from modern semiparametric theory, we obtain a set of influence functions for $\psi_{a^{*}}$, which in turn motivates a feasible set of estimating equations. This step is crucial for developing efficient semiparametric estimators for $\psi_{a^{*}}$ and conducting principled inferences under $\mathcal{M}$.
@@
$\begin{split}\Lambda^{F,\bot}=\bigl{\{}K=&k\left(Y,A,A^{*},Z,X\right):E\left(K\mid Y,A^{*},X\right)\\ &=E\left(K\mid A,A^{*},X\right)=E\left(K\mid Z,A^{\ast},X\right)=0\bigr{\}}\cap L_{0}^{2,F},\end{split}$ (9)
@@
$\begin{split}\Lambda^{F,\bot}=\bigl{\{}W=M-&E\left(M\mid Y,A^{*},X\right)-E\left(M\mid A,A^{*},X\right)-E\left(M\mid Z,A^{*},X\right)\\ &+2E\left(M\mid A^{\ast},X\right):M=m\left(Y,A,A^{*},Z,X\right)\in L_{0}^{2,F}\bigr{\}}.\end{split}$ (10)
The derivation of (9) and (10) is deferred to §LABEL:supp-sec:proof_tangentspace_full in the Supplementary Materials. In the following, we will make reference to the oracle nonparametric model $\mathcal{M}_{np}^{*}$ as the collection of oracle laws compatible with the DAG given in Figure 1(b) which places no restriction on the joint distribution of $(A^{*},Y,Z,A,X)$. We then have the following result:
@@
The EIF of $\psi_{a^{*}}$ in the oracle hidden treatment model $\mathcal{M^{*}}$ is equal to the EIF in the oracle nonparametric model $\mathcal{M}_{np}^{*}$.
@@
The EIF for $\psi_{a^{*}}$ in $\mathcal{M^{*}}_{np}$, denoted by $\phi_{a^{*}}^{F,eff}$ is well-established, and is equal to (2). Therefore, the set of influence functions for $\psi_{a^{*}}$ in the full data model $\mathcal{M^{*}}$ is the linear variety $\phi_{a^{*}}^{F,eff}\oplus\Lambda^{F,\bot}$ (Tsiatis,, 2007).
So far, we have characterized the tangent space, the set of influence functions and the EIF for $\psi_{a^{*}}$ in $\mathcal{M^{*}}$. We now turn to the observed data model $\mathcal{M}$. The following result summarizes our characterization of the tangent space of the observed data hidden treatment model $\Lambda$, and its ortho-complement, obtained by applying a certain coarsening operator which maps the tangent space for the full data model, $\Lambda^{F}$, onto the $L^{2}_{0}(Y,A,Z,X)$ space.
@@
$\begin{split}&\Lambda^{\bot}=\bigl{\{}\{A-E(A\mid A^{*}=0,X)\}\{Z-E(Z\mid A^{*}=1,X)\}h_{1}(Y,X)+\{A-E(A\mid A^{*}=1,X)\}\\ &\{Z-E(Z\mid A^{*}=0,X)\}h_{2}(Y,X):E\{h_{1}(Y,X)\mid A^{*},X\}=E\{h_{2}(Y,X)\mid A^{*},X\}=0\bigr{\}}\cap L^{2}_{0}.\end{split}$ (11)
@@
where $M_{Y\mid A^{*},X}$ is defined as $(v_{1},v_{2})$, with $v_{1}=\left(f(Y=1\mid A^{*}=0,X),\cdots,f(Y=k\mid A^{*}=0,X)\right)^{\mathrm{\scriptscriptstyle T}}$ and $v_{2}=\left(f(Y=1\mid A^{*}=1,X),\cdots,f(Y=k\mid A^{*}=1,X)\right)^{\mathrm{\scriptscriptstyle T}}$. This implies that $h_{1}$ and $h_{2}$ belong to the null space of the $k\times 2$ matrix $M_{Y\mid A^{*},X}$. Under Assumption 2, the null space has a dimension of $k-2$ when $k\geq 2$. When $Y$ is binary, i.e., $k=2$, $M_{Y\mid A^{*},X}$ is full rank, leading to the unique solution of (12): $h_{j}(Y,X)=0,j=1,2$. This is further detailed in Corollary 1.
Corollary 1.
@@
$\Lambda^{\bot}=\bigl{\{}\{A-E(A\mid A^{*}=0,X)\}\{Z-E(Z\mid A^{*}=1,X)\}w_{1}(Y,X)+\{A-E(A\mid A^{*}=1,X)\}$
$\qquad\quad\,\{Z-E(Z\mid A^{*}=0,X)\}w_{2}(Y,X):w_{1},w_{2}\in P_{2}\bigr{\}}.$ (13)
Such equivalence provides a closed-form characterization of $P_{1}$. Based on this, we obtain a closed-form characterization of $\Lambda^{\bot}$. See §LABEL:supp-sec:orthocomplement in the Supplementary Materials for the proof.
@@
The following equation provides a closed-form expression of an influence function for $\psi_{a^{*}}$ in the hidden treatment model $\mathcal{M}$ which applies whether $Y$ is binary, a count or a continuous variable:
@@
$\{Y-E(Y\mid a^{*},X)\}+E(Y\mid a^{*},X)-\psi_{a^{*}},$ (15)
with $Y$ replaced by $\mathbb{I}(Y=y)$ for a categorical $Y$ where $\psi_{a^{*}}=E\{\mathbb{I}(Y^{a^{*}}=y)\}$. Equation (4) admits the triple robustness property: $E(\phi_{a^{*}})=0$ if one of the following holds:
1. $f(A\mid A^{*},X)$ and $f(Y\mid A^{*},X)$ are correctly specified.
2. $f(Z\mid A^{*},X)$ and $f(Y\mid A^{*},X)$ are correctly specified.
3. $f(A\mid A^{*},X)$, $f(Z\mid A^{*},X)$ and $f(A^{*}\mid X)$ are correctly specified.
The EIF $\phi_{a^{*}}^{eff}$ is given by the orthogonal projection of $\phi_{a^{*}}$ onto the observed data tangent space of the hidden treatment model, i.e., $\phi_{a^{*}}^{eff}=\Pi(\phi_{a^{*}}\mid\Lambda)$ where $\Pi$ is the orthogonal projection operator. Therefore,
@@
In the case of binary $Y$, $w^{opt}_{1}=w^{opt}_{2}=0$, and thus (4) is the EIF for $\psi_{a^{*}}$ in the hidden treatment model $\mathcal{M}$. A closed-form expression for $\{w^{opt}_{j}:j=1,2\}$ in the more general categorical outcome setting is given in the Supplemental Materials §LABEL:supp-sec:closed_form_w. As a closed-form expression for $\{w^{opt}_{j}:j=1,2\}$ and thus for the EIF is not generally available for continuous $Y$, we give in the same Supplemental Materials Section a simple approach for obtaining an approximate EIF based on a basis function approximation of $\{w^{opt}_{j}:j=1,2\}$.
The theorem reveals an insightful relationship between an oracle influence function and the corresponding observed data influence function. Specifically, comparing (4) with the full data influence function (2), we observe that the former is obtained by substituting the indicator function $\mathbb{I}_{a^{*}}(A^{*})$ in the latter with:
$M(A,Z,X)=\frac{A-E(A\mid 1-a^{*},X)}{E(A\mid a^{*},X)-E(A\mid 1-a^{*},X)}\frac{Z-E(Z\mid 1-a^{*},X)}{E(Z\mid a^{*},X)-E(Z\mid 1-a^{*},X)}.$ (16)
Interestingly, (16) is a conditionally unbiased estimator of $\mathbb{I}_{a^{*}}(A^{*})$, satisfying $E\{M(A,Z,X)\mid A^{*},X\}=\mathbb{I}_{a^{*}}(A^{*})$. See §LABEL:supp-sec:relationship_if for the derivation. This provides an elegant link between (2) and (4). In (2), the unobserved $A^{*}$ only appears within the indicator function $\mathbb{I}_{a^{*}}(A^{*})$, while in (4), $\mathbb{I}_{a^{*}}(A^{*})$ is replaced by an expression of observed variables (16), whose expectaction conditional on $(A^{*},X)$ equals $\mathbb{I}_{a^{*}}(A^{*})$. This observation suggests a general approach for deriving influence functions that can be extended to other causal functionals of $(A^{*},X,Y)$ discussed in §2 under the hidden treatment model. See §6 for a detailed exposition.